OpenStudy (anonymous):

Factor completely. 5x 2 + 10x - 40 5(x - 4)(x + 2) 5(x - 2)(x + 4) 5(x - 4)(x - 2)

OpenStudy (anonymous):

5(x + 4)(x - 2)

OpenStudy (callisto):

5x^2 + 10x - 40 = 5 (x^2 +2x - 8)  = 5(x+4)(x-2)